# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

### v1

@@ 3.1. Model implication: simulations with a shock
  • mean-field interaction, that corresponds to a symmetric interaction matrix $\Gamma$ of the form
    (3.2) $\gamma_{j,h}=\begin{cases}\gamma^{*}\left(\iota/N+(1-\iota)\right)\quad&\mbox{if }j=h\\ \gamma^{*}\,\iota/N\quad&\mbox{if }j\neq h\end{cases},$
    with $\gamma^{*},\,\iota\in(0,1]$, where we have chosen $\gamma^{*}=0.7$ and $\iota=0.9$, so that
@@
4. Data analysis
Our dataset consists of the GLOBAL PATSTAT database [^8] The dataset is maintained at the IMT School.. More specifically, we collected all the patents published in the period $[1980-2018]$, with their exact (full) date of publication and their CPC-1 category. Moreover, for each patent $n$, we know if it has been cited by subsequent patents (published in the considered period) and which are the citing patents.
[…]
@@ 4. Data analysis
Figure 2. Linear behavior of the successes counter $S_{t,h}$ along $t$, in $\log_{10}-\log_{10}$ scale, for each category $h$. The lines are obtained by a least square interpolation using a subsample of size $10^{4}$ of the data set. We can appreciate how the lines exhibit the same slope, which has been estimated to be equal to $\widehat{\gamma}^{*}=0.689$. Indeed, the goodness of fit $R^{2}$ index obtained imposing a common slope is $0.969$ and so basically equal to the one obtained allowing the slopes to be different across categories, i.e. $0.972$. The emergence of a common Heaps’ exponent across all categories supports the fact that these innovation processes are not evolving independently but are shaped by shared systemic dynamics. Moreover, the sub-linear growth reflects diminishing returns to cumulative knowledge, increasing innovation complexity, and a bounded scalability in innovation outputs over time.
[…]
@@
Figure 3. Plot of the process $S_{t,h}/S_{t,H}$ along time $t$, for all the categories $h\neq H$. Category $H$ is arbitrarily chosen to be the baseline category. The horizontal dashed red lines represent an estimation of the ratio $u_{h}/u_{H}$, obtained as $10^{d_{h,H}}$, where $d_{h,H}$ is the difference between the intercepts of the regression lines in Figure 2 for the categories $h$ and $H$. This long-run convergence across categories highlights a structural stabilization in innovation outcomes, despite increasing complexity and declining marginal returns to knowledge production. Moreover, note that all the estimated ratios $u_{h}/u_{H}$ are around $1$. This means that in the long run the numbers of successes in the different categories tend to coincide.
@@
Figure 3 shows, for each category $h$, the convergence of the process $(S_{t,h}/S_{t,H})_{t}$ toward the quantity $u_{h}/u_{H}$, estimated as $10^{d_{h,H}}$ where $d_{h,H}$ is the difference between the intercepts of the regression lines in Figure 2 for the two processes related to the pair $(h,H)$ of categories. [^11] Category H in Figure 3 is arbitrarily chosen to be the baseline category. Any other category can be used as a baseline. Also this fact is in accordance with Theorem 3.3 and it means that the relative number of successes across categories stabilizes along time. This effect reflects the fact that, in the long run, the rates at which capabilities evolve into core capabilities—and core capabilities into rigidities—within a technological domain, including the cross-sectoral influences, tend to settle. In other words, the temporal stabilization of innovation successes implies two distinct mechanisms: (i) the convergence of the catalytic effect of general-purpose technologies on innovation across diverse domains (Trajtenberg et al. 1997), as reflected in the equilibrium configuration of the subset of $\Gamma$ related to such technologies; and (ii) the stabilization in the rate of recombination of existing knowledge, as captured by the behavior of $P_{t,h}$ over time. Moreover, the limit corresponds to the ratio of their respective components of the eigenvector-centrality scores ${\mathbf{u}}$ (with respect to $\Gamma^{\top}$). In Figure 3 we can note that the estimated quantities $u_{h}/u_{H}$ are all very near to $1$ (more precisely, they are all within the interval $[1,1.3]$). This means that in the long run the numbers of successes in the different categories tend to coincide, which carries significant implications for innovation policy. In the case of the green transition, for instance, it means that even if different technology areas seem to be producing similar results over time, this similarity can hide the fact that it’s actually getting harder and more complex to create those innovations (Bloom et al. 2020).
@@
Figure 4. Linear behavior of the successes counter $S_{t,k,h}$ along $t$ in $\log_{10}-\log_{10}$ scale, for each pair $(k,h)$ of categories ($k=$sub-figure, $h=$color). The lines are obtained by a least square interpolation based on a subsample of the dataset with a slope equal to the previously estimated value $\widehat{\gamma}^{*}=0.689$. The goodness of fit $R^{2}$ index is $0.851$, which is basically the same as the one obtained allowing the slopes to be different across the pairs $(k,h)$ of categories, i.e. $0.855$. The common slope supports one again the presence of interaction across the categories. Moreover, the sub-linear growth reflects a consistent sub-linear scaling of cross-domain innovation flows, suggesting that while knowledge recombination persists, its marginal productivity diminishes over time across all category pairs.
@@
Figure 5. Plot of the process $S_{t,k,h}/S_{t,H,h}$ along time $t$, for all the pairs $(k,h)$ of categories ($k=$sub-figure, $h=$color), with $k\neq H$. Category $H$ is arbitrarily chosen to be the baseline category. The horizontal dashed red lines represent the value $10$ to the power of the differences between the intercepts of the regression lines in Figure 4 for the pairs of categories $(k,h)$ and $(H,h)$. This long-run proportionality links early patenting intensity to long-term innovation success, highlighting the path-dependent and self-reinforcing nature of technological trajectories.
@@
4.1. Statistical inference on the interaction intensity under the mean-field assumption
Assuming the system is subjected to a mean-field interaction (note that in this case we have $u_{h}/u_{j}=1$ for each $h,j\in{\mathcal{H}}$), i.e. assuming the interaction matrix $\Gamma$ to be of the form (3.2), we can perform a suitable test on the parameter $\iota$, that rules the intensity of the interaction. Indeed, since the leading eigenvalue $\gamma^{*}$ can be estimated with high accuracy by a suitable regression on the vector process ${\mathbf{S}}_{t}=(S_{t,h})_{h\in{\mathcal{H}}}^{\top}$ (see Aletti et al. (2023a) and references therein), we can assume this parameter to be known and we can make inference only on the intensity parameter $\iota$. For a two-sided test with $H_{0}:\;\iota=\iota_{0}$, where $\iota_{0}\leq 1$ and $\iota_{0}>1/2$ (the required condition in the theoretical result for having the second eigenvalue of $\Gamma$ strictly smaller than $\gamma^{*}/2$), we can use the test statistic (see Appendix B for the technical details)
(4.1) $2\Delta_{0}\frac{\big{\|}{\mathbf{S}}_{t}-\widetilde{S}_{t}{\boldsymbol{1}}\big{\|}^{2}}{\widetilde{S}_{t}}\stackrel{{\scriptstyle d}}{{\sim}}\chi^{2}(N-1)\quad\mbox{under }H_{0},$
where $\Delta_{0}=\iota_{0}-\frac{1}{2}$ and $\widetilde{S}_{t}=\tfrac{{\boldsymbol{1}}^{\top}{\mathbf{S}}_{t}}{N}=\sum_{h\in{\mathcal{H}}}S_{t,h}/N$. Since this statistic is increasing in $\iota_{0}$, it works well also for the one-sided test with $H_{0}:\;\iota\geq\iota_{0}$, with $\iota_{0}\in(1/2,1]$. Applying this one-sided test to our real dataset of patents (with the estimated value $\widehat{\gamma}^{*}=0.689$), we obtain different $p$-values, one for each tested value $\iota_{0}$, that are collected in Table 1. We can see that, if we choose the significance level $\alpha=0.05$, then the minimum value of $\iota_{0}$ at which we can reject the null hypothesis is $\iota_{0}=0.75$.
We have also computed an estimate of the parameter of interest $\iota$ maximizing the (joint) likelihood of the processes $X^{(h)}=(X_{n,h})_{n}$, $h\in{\mathcal{H}}$, i.e.
@@ 4.1. Statistical inference on the interaction intensity under the mean-field assumption
p-value 0.893 0.562 0.273 0.113 0.042 0.015 0.005 0.002 $<0.001$
Table 1. $p$-values associated to the hypothesis test with $H_{0}:\iota\geq\iota_{0}$ under the mean-field interaction .
Our estimate $\hat{\iota}$ of the intensity parameter $\iota$ shows that on the one hand, processes are interdependent, i.e. they share knowledge, influence, or innovation spillovers and on the other, that self-dynamics still matter. This reinforces the concept of dynamic competitive advantage as introduced by Teece et al. (1997) and Giannitsis and Kager (2009). Broadly speaking, the transformation of general capabilities into core capabilities occurs through a process of smart specialization. In this process, technological capabilities must adapt to multidisciplinary stimuli originating from other technologies, translate them into their own technological language, and leverage them to explore new opportunities within their own domain.
5. Conclusions
@@
Beyond its theoretical contributions, the framework offers practical relevance for managers. It namely underscores the importance of early and well-targeted interventions—particularly before ecosystems fragment into highly specialized, path-dependent trajectories (David 1985, Mazzucato 2018). The model further suggests that supporting cross-domain connectivity (e.g., via interdisciplinary R&D programs or incentives for recombination across patent classes) may delay the onset of technological rigidities and foster broader capability evolution (Jacobides et al. 2018, Kim and Magee 2017). Finally, by identifying the statistical structure of innovation interactions, our model provides policymakers with a novel tool to detect when technological fields are becoming saturated—informing decisions on resource allocation and diversification strategies in national or regional innovation systems (Fleming 2001, Jaffe and Trajtenberg 2002).
@@
Author contributions
All the authors contributed equally to the present work.
Acknowledgments
Fundings
Irene Crimaldi and Federico Nutarelli are partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research under the National Recovery and Resilience Plan (PNRR) - M4C2 - Investment 1.3, title [Public Sector Indicators for Sustainability and Wellbeing (PUWELL)] - Program [Growing Resilient, INclusive and Sustainable (GRINS)] - PE18 - CUP J33C22002910001.
This research was supported by the “Resilienza Economica e Digitale” (RED) project
(CUP D67G23000060001) funded by the Italian Ministry of University and Research (MUR) as “Department of Excellence” (Dipartimenti di Eccellenza 2023-2027, Ministerial Decree no. 230/2022)
Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes. Supplementary material
@@ Appendix A Definitions of novelty in the literature
For a patent to be defined as such, it must be “novel” at the time of its filing (Mueller 2024). In other words, a patent is novel by definition, and the concept of novelty is widely discussed in the literature (Abbas et al. 2014). Is novelty related to how impactful a patent is? Or how influential it becomes? Or perhaps how unexpected its release was?
@@
Squicciarini et al. (2013) $\text{Originality}_{p}=1-\sum_{j}^{n_{p}}s_{pj}^{2}$ $s_{pj}$ is the percentage of citations made by patent p to patent class $j$ out of the $n_{p}$ IPC 4-digit (or 7-digit) patent codes contained in the patents cited by patent $p$. Citation measures are built on EPO patents and account for patent equivalents. Citations
Squicciarini et al. (2013) $\text{Radicalness}_{p}=\sum_{j}^{n_{p}}\frac{\text{CT}_{j}}{n_{p}}\,;\,\text{IPC}_{pj}\neq\text{IPC}_{p}$ $CT_{j}$ denotes the count of IPC-4 digit codes $CPC_{pj}$ of patent $j$ cited in patent $p$ that is not allocated to patent $p$, out of $n$ CPC classes in the backward citations counted at the most disaggregated level available (up to the 5th hierarchical level). The higher the ratio, the more diversified the array of technologies on which the patent relies upon. Citations Tech.Knowledge
Own elaboration based on intuitions in Hall et al. (2000) and Lanjouw and Schankerman (2004) $\text{ANPCI}_{i,t}=\frac{C_{i,t}}{\lambda_{t}\cdot\left(\frac{1}{n_{t}}\sum_{j=1}^{n_{t}}C_{j,t}\right)}$ where: - $C_{i,t}$ is the number of citations received by patent $i$ in year $t$, - $n_{t}$ is the number of patents filed in year $t$, - $\frac{1}{n_{t}}\sum_{j=1}^{n_{t}}C_{j,t}$ is the average number of citations per patent in year $t$, -$\lambda_{t}$ is a correction factor to account for citation inflation, which adjusts the average citation counts for inter-year comparability. Is the average number of citations per patent across all fields in year $t$, divided by a baseline year (e.g., $t_{0}$) Intra-year comparison: index is normalized by the average citations within the same cohort (year and field). Inter-year comparison: adjusting for citation inflation $\lambda_{t}$ with , we should mitigate the bias of comparing older patents with more citations to newer ones that haven’t had as much time to accumulate citations. Citations
Trajtenberg et al. (1997) (Modified better version) $G_{X}=1-\sum_{j=1}^{M_{i}}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{T_{ji}^{n}}{T_{i}^{n}}\right)^{2}$ $X$ is the focal patent with $Y_{i}$ patents citing the focal patent $X$, with $i=1,\cdot,N$. $T_{i}^{n}$ is the total number of IPC $n$-digit classes in $y_{i}$ $T_{ji}^{n}$ is the total number of IPC $n$-digit classes in the $j^{th}$ IPC4 digit class in $y_{i}$ and $j=1\dots M_{i}$ is the cardinal of all IPC4-digit classes in $y_{i}$ Citations Tech.Knowledge
Various Patent quality: Composite indices Typically based on patent citations, claims, patent renewals and patent family size. Usually different compositions as follows: [i]Patent quality index 4 – 4 components: number of forward citations (up to 5 years after publication); patent family size; number of claims; and the patent generality index. Only granted patents are covered by the index. [ii] Patent quality index 4b – 4 components, bis: number of forward citations (up to 5 years after publication); patent family size; corrected claims; and the patent generality index. Only granted patents are covered by the index. [iii] Patent quality index 6 – 6 components: covers the same components as above, plus the number of backward citations and the grant lag index. Entropy based approaches: Defines a weighting scheme ($W$) for all indicators of a patent (citations, claims…) based on entropy. With $M$ patents, it then constructs $M\cdot W$. A set of negative (patents with the lowest weighted scores across the majority of indicators) and positive patents is then identified and divided; a similarity across good and negative patents is then computed. The following steps are then performed: Maximum Similarity: For each patent in $M^{R}$, the maximum similarity to any patent in $M^{N}$ is determined. If the maximum similarity $S(M_{i}^{R},M^{N})$ exceeds a predefined threshold $\tau$, that remaining patent is marked as negative. The process repeats iteratively. Newly marked negative patents are moved to $M^{N}$, and the model re-evaluates the remaining patents in $M^{R}$. The iteration stops when the number of remaining patents $M^{R}$ is less than a threshold $\theta$, or no new patents are marked as negative. The final output is a small set of patents in $M^{R}$ that are not marked as negative, indicating their higher potential for technological innovation based on distinctiveness and low similarity to the negative patents. Various
@@
Table 2. This Table lists the various novelty indices brought by the literature. As noted in Lanjouw and Schankerman (2004), however, no single indicator can fully capture the quality or value of a patent and multiple indicators (citations, claims, oppositions, family size) provide a more comprehensive and nuanced understanding of both the technological importance and the economic potential of innovations.
@@
    The older and more established literature is the one that defines novelty basing on the forward and backward citations (Trajtenberg et al. 1997, Squicciarini et al. 2013). Moreover, in the document about PATSTAT database we can read: “The number of citing patent documents can be an indicator of the importance of the patent. A frequently cited patent can be an indication of a core technology”. This literature have been recently criticized. In Abbas et al. (2014), for instance, the authors primarily discuss patent novelty in the context of citation analysis, classification systems, and keywords. They highlight how backward citations (patents cited by a new patent) are commonly used to infer novelty—specifically, fewer backward citations can signal higher novelty because the patent draws less from existing technologies. However, they criticize citation-based methods, noting that they don’t always capture the full semantic or functional novelty of a patent.

### v2

@@ 3.1 Model implication: simulations with a shock
  • mean-field interaction, that is a symmetric interaction matrix $\Gamma$ of the form
    $\gamma_{j,h}=\begin{cases}\gamma^{*}\left(\iota/N+(1-\iota)\right)\quad&\mbox{if }j=h\\ \gamma^{*}\,\iota/N\quad&\mbox{if }j\neq h\end{cases},$ (2)
    with $\gamma^{*},\,\iota\in(0,1]$  [^5] We refer to this kind of interaction as the mean-field interaction, because in this case the term $\sum_{j\in\mathcal{H}}\gamma_{j,h}S_{t,j}$ in (1) becomes $\gamma^{*}\left(\iota\sum_{j\in{\mathcal{H}}}S_{t,j}/N+(1-\iota)S_{t,h}\right)$ and so the conditional probability $P_{t,h}$ depends on a convex combination between $S_{t,h}$ and the averaged value of all the $S_{t,j}$ in the system., where we have chosen $\gamma^{*}=0.7$ and $\iota=0.9$, so that
@@
4 Data analysis
Our dataset consists of the GLOBAL PATSTAT database [^6] The dataset is maintained at the IMT School.. More specifically, we collected all the patents published in the period [19802018][1980-2018], with their exact (full) date of publication and their CPC-1 category. Moreover, for each patent $n$, we know if it has been cited by subsequent patents (published in the considered period) and which are the citing patents.
We consider the patents with publication year in the period 198020131980-2013 so that, for each of them, we are able to compute the index $I_{n,h}$. We limit ourselves to the CPC categories $h\in{\mathcal{H}}=\{A,\,B,\,C,\,D,\,E,\,F,\,G,\,H\}$ (i.e. we exclude category $Y$  [^7] This category has been excluded due to its broad scope and ambiguity, which are well-documented in the literature (Leydesdorff, Kogler and Yan 2017, Rainville, Dikker and Buggenhagen 2025). Although it is often broadly referred to as the class encompassing green patents, it is widely recognized in the green innovation literature that only the subclasses $Y02$ and $Y04$ accurately represent genuine green technologies (Corrocher and Mancusi 2021, Barbieri, Consoli and Napolitano 2023).). We fix $T=5$ and the threshold $\tau=0.8$ (see Section D for some details on the choice of the threshold). We thus obtain a matrix with $N=\text{card}({\mathcal{H}})=8$ columns, where each row corresponds to a patent $n$. The patents (and so the rows) are ordered with respect to their publication (full) date. The total number of rows is $n_{tot}=5\,004\,253$. The matrix entry $x_{n,h}$ is equal to $1$ if the patent $n$ is a success for category $h$, i.e. if its value of the index is above the threshold $\tau$, or equal to 0 if $n$ is not a success. Hence, the constructed matrix can be seen as the realization of a finite system $\{X^{(h)}=(X_{n,h})_{n}\,:\,h\in{\mathcal{H}}\}$ of Bernoulli processes. Our analyses show that the real data $\{x_{n,h}\}$ exhibit a behavior along time in agreement with the theoretical results of the previous section.
@@ 4 Data analysis
Figure 2: Linear behavior of the successes counter $S_{t,h}$ along $t$, in $\log_{10}-\log_{10}$ scale, for each category $h$. The lines are obtained by a least square interpolation using a subsample of size $10^{4}$ of the data set. We can appreciate how the lines exhibit the same slope, which has been estimated to be equal to $\widehat{\gamma}^{*}=0.689$. Indeed, the goodness of fit $R^{2}$ index obtained imposing a common slope is $0.969$ and so basically equal to the one obtained allowing the slopes to be different across categories, i.e. $0.972$. The emergence of a common Heaps’ exponent across all categories supports the fact that these innovation processes are not evolving independently but are shaped by shared systemic dynamics. Moreover, the sub-linear growth reflects diminishing returns to cumulative knowledge, increasing innovation complexity, and a bounded scalability in innovation outputs over time.
The value of the common Heaps’ exponent, estimated as the common slope of the lines in the $\log_{10}-\log_{10}$ plot, is $\widehat{\gamma}^{*}=0.689$. [^8] In order to measure the uncertainty of the slopes estimated for each category $h$ at a given time-step, we have performed a linear regression with random effects, on both intercept and slope and using a subsample of size $10^{4}$ of the data set. The estimated values of the slopes, one for each category $h$, presents a mean equal to $\widehat{\gamma}^{*}=0.689$ and standard deviation equal to $0.044$. Therefore, the number of successes in each category grows sub-linearly, i.e. we have a decreasing (as $t^{-(1-\widehat{\gamma}^{*})}$) probability of having a success and vanishing (as $t^{-(1-\widehat{\gamma}^{*})}$) correlation coefficients among the categories, mirroring increasing complexity, higher search costs, and diminishing returns to cumulative knowledge in each domain. This uniform sub-linear growth pattern across categories reflects a form of bounded scalability in innovation outputs: while innovation continues over time, the marginal probability of producing a new success in any category diminishes. This aligns with the recently emerging notion that innovation becomes harder as technological frontiers advance (Bloom et al. 2020).
@@
Figure 3: Plot of the process $S_{t,h}/S_{t,H}$ along time $t$, for all the categories $h\neq H$. Category $H$ is arbitrarily chosen to be the baseline category. The horizontal dashed red lines represent an estimation of the ratio $u_{h}/u_{H}$, obtained as $10^{d_{h,H}}$, where $d_{h,H}$ is the difference between the intercepts of the regression lines in Figure 2 for the categories $h$ and $H$. This long-run convergence across categories highlights a structural stabilization in innovation outcomes, despite increasing complexity and declining marginal returns to knowledge production. Moreover, note that all the estimated ratios $u_{h}/u_{H}$ are around 11. This means that in the long run the numbers of successes in the different categories tend to coincide.
@@
Figure 3 shows, for each category $h$, the convergence of the process $(S_{t,h}/S_{t,H})_{t}$ toward the quantity $u_{h}/u_{H}$, estimated as $10^{d_{h,H}}$ where $d_{h,H}$ is the difference between the intercepts of the regression lines in Figure 2 for the two processes related to the pair $(h,H)$ of categories. [^9] Category H in Figure 3 is arbitrarily chosen to be the baseline category. Any other category can be used as a baseline. Also this fact is in accordance with Theorem 3.2 and it means that the relative number of successes across categories stabilizes along time. This effect reflects the fact that, in the long run, the rates at which capabilities evolve into core capabilities—and core capabilities into rigidities—within a technological domain, including the cross-sectoral influences, tend to settle. In other words, the temporal stabilization of innovation successes implies two distinct mechanisms: (i) the convergence of the catalytic effect of general-purpose technologies on innovation across diverse domains (Trajtenberg, Henderson and Jaffe 1997), as reflected in the equilibrium configuration of the subset of $\Gamma$ related to such technologies; and (ii) the stabilization in the rate of recombination of existing knowledge, as captured by the behavior of $P_{t,h}$ over time. Moreover, the limit corresponds to the ratio of their respective components of the eigenvector-centrality scores $\mathbf{u}$} (with respect to $\Gamma^{\top}$). In Figure 3 we can note that the estimated quantities $u_{h}/u_{H}$ are all very near to $1$ (more precisely, they are all within the interval $[1,1.3]$). This means that in the long run the numbers of successes in the different categories tend to coincide, which carries significant implications for innovation policy. In the case of the green transition, for instance, it means that even if different technology areas seem to be producing similar results over time, this similarity can hide the fact that it’s actually getting harder and more complex to create those innovations (Bloom et al. 2020).
@@
Figure 4: Linear behavior of the successes counter $S_{t,k,h}$ along $t$ in $\log_{10}-\log_{10}$ scale, for each pair $(k,h)$ of categories (k=k=sub-figure, h=h=color). The lines are obtained by a least square interpolation based on a subsample of the dataset with a slope equal to the previously estimated value $\widehat{\gamma}^{*}=0.689$. The goodness of fit $R^{2}$ index is $0.851$, which is basically the same as the one obtained allowing the slopes to be different across the pairs $(k,h)$ of categories, i.e. $0.855$. The common slope supports one again the presence of interaction across the categories. Moreover, the sub-linear growth reflects a consistent sub-linear scaling of cross-domain innovation flows, suggesting that while knowledge recombination persists, its marginal productivity diminishes over time across all category pairs.
@@
Figure 5: Plot of the process $S_{t,k,h}/S_{t,H,h}$ along time $t$, for all the pairs $(k,h)$ of categories (k=k=sub-figure, h=h=color), with $k\neq H$. Category $H$ is arbitrarily chosen to be the baseline category. The horizontal dashed red lines represent the value $10$ to the power of the differences between the intercepts of the regression lines in Figure 4 for the pairs of categories $(k,h)$ and $(H,h)$. This long-run proportionality links early patenting intensity to long-term innovation success, highlighting the path-dependent and self-reinforcing nature of technological trajectories.
@@
4.1 Statistical inference on the interaction intensity under the mean-field assumption
Assuming the system is subjected to a mean-field interaction (note that in this case we have $u_{h}/u_{j}=1$ for each $h,j\in{\mathcal{H}}$), i.e. assuming the interaction matrix $\Gamma$ to be of the form (2), we can perform a suitable test on the parameter $\iota$, that rules the intensity of the interaction. For a two-sided test with $H_{0}:\;\iota=\iota_{0}$, where $\iota_{0}\leq 1$ and $\iota_{0}>1/2$ (the required condition in the theoretical result for having the second eigenvalue of $\Gamma$ strictly smaller than $\gamma^{*}/2$), we can use the test statistic (see Appendix B for the technical details)
$2\Delta_{0}\frac{\big\|{\mathbf{S}}_{t}-\widetilde{S}_{t}{\mathbf{1}}\big\|^{2}}{\widetilde{S}_{t}}\stackrel{{\scriptstyle d}}{{\sim}}\chi^{2}(N-1)\quad\mbox{under }H_{0},$ (3)
where $\Delta_{0}=\iota_{0}-\frac{1}{2}$ and $\widetilde{S}_{t}=\tfrac{{\mathbf{1}}^{\top}{\mathbf{S}}_{t}}{N}=\sum_{h\in{\mathcal{H}}}S_{t,h}/N$. Since this statistic is increasing in $\iota_{0}$, it works well also for the one-sided test with $H_{0}:\;\iota\geq\iota_{0}$, with $\iota_{0}\in(1/2,1]$. Applying this one-sided test to our real dataset of patents, we obtain different $p$-values, one for each tested value $\iota_{0}$, that are collected in Table 1. We can see that, if we choose the significance level $\alpha=0.05$, then the minimum value of $\iota_{0}$ at which we can reject the null hypothesis is $\iota_{0}=0.75$.
Taking the value of the parameter $\gamma^{*}$ equal to the estimated one, i.e. $\widehat{\gamma}^{*}=0.689$, we have also computed an estimate of the parameter of interest $\iota$ maximizing the (joint) likelihood of the processes $X^{(h)}=(X_{n,h})_{n}$, $h\in{\mathcal{H}}$, i.e.
@@ 4.1 Statistical inference on the interaction intensity under the mean-field assumption
p-value 0.893 0.562 0.273 0.113 0.042 0.015 0.005 0.002 <0.001<0.001
Table 1: $p$-values associated to the hypothesis test with $H_{0}:\iota\geq\iota_{0}$ under the mean-field interaction .
Our estimate $\hat{\iota}$ of the intensity parameter $\iota$ shows that on the one hand, processes are interdependent, i.e. they share knowledge, influence, or innovation spillovers and on the other, that self-dynamics still matter. This reinforces the concept of dynamic competitive advantage as introduced by Teece, Pisano and Shuen (1997) and Giannitsis and Kager (2009). Broadly speaking, the transformation of general capabilities into core capabilities occurs through a process of smart specialization. In this process, technological capabilities must adapt to multidisciplinary stimuli originating from other technologies, translate them into their own technological language, and leverage them to explore new opportunities within their own domain.
5 Conclusions
@@
Beyond its theoretical contributions, the framework offers practical relevance for managers. It namely underscores the importance of early and well-targeted interventions—particularly before ecosystems fragment into highly specialized, path-dependent trajectories (David 1985, Mazzucato 2018). The model further suggests that supporting cross-domain connectivity (e.g., via interdisciplinary R&D programs or incentives for recombination across patent classes) may delay the onset of technological rigidities and foster broader capability evolution (Jacobides, Cennamo and Gawer 2018, Kim and Magee 2017). Finally, by identifying the statistical structure of innovation interactions, our model provides policymakers with a novel tool to detect when technological fields are becoming saturated—informing decisions on resource allocation and diversification strategies in national or regional innovation systems (Fleming 2001, Jaffe and Trajtenberg 2002).
@@ Appendix A Definitions of novelty in the literature
For a patent to be defined as such, it must be “novel” at the time of its filing (Mueller 2024). In other words, a patent is novel by definition, and the concept of novelty is widely discussed in the literature (Abbas, Zhang and Khan 2014). Is novelty related to how impactful a patent is? Or how influential it becomes? Or perhaps how unexpected its release was?
@@
Squicciarini, Dernis and Criscuolo (2013) $\text{Originality}_{p}=1-\sum_{j}^{n_{p}}s_{pj}^{2}$
$s_{pj}$ is the percentage of citations made by patent p to patent class $j$ out of the
$n_{p}$ IPC 4-digit (or 7-digit) patent codes contained in the patents cited by patent $p$.
Citation measures are built on EPO patents and account for patent equivalents.
Citations
Squicciarini, Dernis and Criscuolo (2013) $\text{Radicalness}_{p}=\sum_{j}^{n_{p}}\frac{\text{CT}_{j}}{n_{p}}\,;\,\text{IPC}_{pj}\neq\text{IPC}_{p}$
$CT_{j}$ denotes the count of IPC-4 digit codes $CPC_{pj}$ of patent $j$ cited in patent $p$ that is
not allocated to patent $p$, out of $n$ CPC classes in the backward citations counted at
the most disaggregated level available (up to the 5th hierarchical level). The higher
the ratio, the more diversified the array of technologies on which the patent relies upon.
Citations
Tech.Knowledge
Own elaboration based
on intuitions in Hall, Jaffe and Trajtenberg (2000)
and
Lanjouw and Schankerman (2004)
$\text{ANPCI}_{i,t}=\frac{C_{i,t}}{\lambda_{t}\cdot\left(\frac{1}{n_{t}}\sum_{j=1}^{n_{t}}C_{j,t}\right)}$
where: - $C_{i,t}$ is the number of citations
received by patent $i$ in year $t$,
- $n_{t}$ is the number of patents
filed in year $t$,
- $\frac{1}{n_{t}}\sum_{j=1}^{n_{t}}C_{j,t}$ is the
average number of citations per patent in year $t$,
-$\lambda_{t}$ is a correction factor to account for
citation inflation, which adjusts the
average citation counts for inter-year comparability.
Is the average number of citations per patent across
all fields in year $t$, divided by a baseline year (e.g., $t_{0}$)
Intra-year comparison: index is normalized by the average
citations within the same cohort (year and field).
Inter-year comparison: adjusting for citation inflation $\lambda_{t}$ with
, we should mitigate the bias of comparing
older patents with more citations to newer
ones that haven’t had as much time to accumulate citations.
Citations
Trajtenberg, Henderson and Jaffe (1997)
(Modified better version)
$G_{X}=1-\sum_{j=1}^{M_{i}}\left(\frac{1}{N}\sum_{i=1}^{N}\frac{T_{ji}^{n}}{T_{i}^{n}}\right)^{2}$
$X$ is the focal patent with $Y_{i}$ patents citing the focal patent $X$, with $i=1,\cdot,N$.
$T_{i}^{n}$ is the total number of IPC $n$-digit classes in $y_{i}$
$T_{ji}^{n}$ is the total number of IPC $n$-digit classes in the $j^{th}$ IPC4 digit class in $y_{i}$
and $j=1\dots M_{i}$ is the cardinal of all IPC4-digit classes in $y_{i}$
Citations
Tech.Knowledge
Various
Patent quality:
Composite indices
Typically based on patent citations, claims, patent renewals and patent family size.
Usually different compositions as follows: [i]Patent quality index 4 – 4 components:
number of forward citations (up to 5 years after publication); patent family size;
number of claims; and the patent generality index. Only granted patents are covered
by the index. [$i$] Patent quality index 4b – 4 components, bis: number of
forward citations (up to 5 years after publication); patent family size; corrected claims; and
the patent generality index. Only granted patents are covered by the index.
[iii] Patent quality index 6 – 6 components: covers the same components
as above, plus the number of backward citations and the grant lag index.
Entropy based approaches: Defines a weighting scheme ($W$) for all indicators of a patent (citations, claims…)
based on entropy. With $M$ patents, it then constructs $M\cdot W$. A set of negative
(patents with the lowest weighted scores across the majority of indicators) and positive patents is
then identified and divided; a similarity across good and negative patents is then computed.
The following steps are then performed: Maximum Similarity: For each patent in
$M^{R}$, the maximum similarity to any patent in $M^{N}$ is determined. If the maximum similarity $S(M_{i}^{R},M^{N})$ exceeds
a predefined threshold $\tau$, that remaining patent is marked as negative.
The process repeats iteratively. Newly marked negative patents are moved to $M^{N}$, and the model
re-evaluates the remaining patents in $M^{R}$. The iteration stops when the number of remaining
patents $M^{R}$ is less than a threshold $\theta$, or no new patents are marked as negative.
The final output is a small set of patents in $M^{R}$ that are
not marked as negative, indicating their higher potential for technological innovation based
on distinctiveness and low similarity to the negative patents.
Various
@@
Table 2: This Table lists the various novelty indices brought by the literature. As noted in Lanjouw and Schankerman (2004), however, no single indicator can fully capture the quality or value of a patent and multiple indicators (citations, claims, oppositions, family size) provide a more comprehensive and nuanced understanding of both the technological importance and the economic potential of innovations.
@@
    The older and more established literature is the one that defines novelty basing on the forward and backward citations (Trajtenberg, Henderson and Jaffe 1997, Squicciarini, Dernis and Criscuolo 2013). Moreover, in the document about PATSTAT database we can read: “The number of citing patent documents can be an indicator of the importance of the patent. A frequently cited patent can be an indication of a core technology”. This literature have been recently criticized. In Abbas, Zhang and Khan (2014), for instance, the authors primarily discuss patent novelty in the context of citation analysis, classification systems, and keywords. They highlight how backward citations (patents cited by a new patent) are commonly used to infer novelty—specifically, fewer backward citations can signal higher novelty because the patent draws less from existing technologies. However, they criticize citation-based methods, noting that they don’t always capture the full semantic or functional novelty of a patent.